40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 6858     Entry time: Fri Jun 22 20:58:15 2012
Author: Jenne 
Type: Update 
Category: Green Locking 
Subject: Calibrated POX spectra - Xarm stabilized by ALS 

[Yuta, Jenne, Koji]

We stabilized the Xarm using the ALS and took a spectrum of POX as our out of loop sensor.  We used the calibration from elog 6841 to go from counts to meters.

We find (see attached pdf) that the RMS is around 60pm, dominated by 1Hz motion.

 

In other, related, news, I took out the beam pipe connecting the AP and PSL tables and covered the holes with foil.  This makes it much easier and faster to get to the X beat setup for alignment.  Eventually we'll have to put it back, but while the AUX laser on the AP table is not being used for beating against the PSL it'll be nice to have it out of the way.

Attachment 1: POX11_I_ERR_calib_residualCavityMotion_better.pdf  21 kB  | Hide | Hide all
POX11_I_ERR_calib_residualCavityMotion_better.pdf
ELOG V3.1.3-